OpenStudy (jadedry):

Sin 3x + sin 6x + sin 9x = 0 - Solve, giving values from 0 -360. (Half completed need help with second part please!)

OpenStudy (jadedry):

First off i converted it to: \[2\sin6x \cos3x +\sin 6x = 0\] Therefore, Sin 6x or Cos 3x = -0.5 . for Sin 6X, x = \[\frac{ n \pi }{ 6 }\] = 0,30,60,90,120....360 For Cos 3X to make the entire equation 0 = 3x = 120, x = 40 But from this point forward i'm a bit confused as to how to proceed? I thought I would have values of 40, 120, 200. But apparently not? Thanks in advance!

OpenStudy (sirm3d):

\[2\sin6x\cos3x+\sin6x=0\] \[\sin6x\left\{2\cos3x+1\right\}=0\] \[\sin6x=0,2\cos3x+1=0\] solve the two equations separately

OpenStudy (sirm3d):

also, when solving \(\sin6x=0\) for \(0\le x<360\), you should consider the range of values \(0\le 6x <6\times360\)

OpenStudy (jadedry):

I had those equations at one point, I'm still confused about cos though, if cos 3x + 1 = 0 then x = 60 which is wrong according to my book? Maybe I'm misunderstanding you?

OpenStudy (sirm3d):

\[2\cos3x+1=0\] \[\cos3x=-\frac{1}{2}\] \[3x=120,240,480,600,840,960\] therefore \[x=40,80,160,200,280,320\]

OpenStudy (jadedry):

Ah! I understand. I was dividing it before I came up with a list of numbers. Thank you. c:

OpenStudy (sirm3d):

when \(\sin6x=0\) then \[6x=0,180,360,540, 720,900,1080,1260,1440,1620,1800,1980\] or \[x=0,300,60,90, 120,150,180,210,240,270,300,330\]

OpenStudy (sirm3d):

you're welcome

OpenStudy (jadedry):

Alright, gotcha, thanks again! c: